Determine if the following are solutions to the inequality...
-1, 1, 6, 13
x - 3>_6 (the _ is supposed to be inderneath the >)
Can someone please help me, I do not understand how to do these at all.

Answer by oscargut(2103)

-1, 1, 6, 13
x - 3>_6
you have to substitute values and see if the inequality verifies:
if x=-1 then -1-3>_6 then -4>_6 (then x=-1 is not a solution to the inequality)
if x= 1 then 1-3>_6 then -2>_6 (then x=1 is not a solution to the inequality)
if x= 6 then 6-3>_6 then 3>_6 (then x=6 is not a solution to the inequality)
if x=13 then 13-3>_6 then 10>_6 (then x=13 is a solution to the inequality)